DETERMINE HOW MUCH LONGER A 1.00S PROPER TIME INTERVAL APPEARS TO A STATIONARY OBSERVER WHEN A CLOCK IS MOVING WITH A SPEED OF 0.6C


1
Expert's answer
2021-05-03T10:33:59-0400
"\Delta t=\frac{t}{\sqrt{1-\frac{v^2}{c^2}}}-t\\
\Delta t=\frac{1}{\sqrt{1-0.6^2}}-1=0.25\ s"